# KG-CQR: Leveraging Structured Relation Representations in Knowledge Graphs for Contextual Query Retrieval

### v2

@@ Abstract
The integration of knowledge graphs (KGs) with large language models (LLMs) offers significant potential to improve the retrieval phase of retrieval-augmented generation (RAG) systems. In this study, we propose KG-CQR [^1] https://github.com/tnmai59/KG-CQR, a novel framework for Contextual Query Retrieval (CQR) that enhances the retrieval phase by enriching the contextual representation of complex input queries using a corpus-centric KG. Unlike existing methods that primarily address corpus-level context loss, KG-CQR focuses on query enrichment through structured relation representations, extracting and completing relevant KG subgraphs to generate semantically rich query contexts. Comprising subgraph extraction, completion, and contextual generation modules, KG-CQR operates as a model-agnostic pipeline, ensuring scalability across LLMs of varying sizes without additional training. Experimental results on the RAGBench and MultiHop-RAG datasets demonstrate that KG-CQR outperforms strong baselines, achieving improvements of up to 4–6% in mAP and around 2–3% in Recall@25. Furthermore, evaluations on challenging RAG tasks such as multi-hop question answering show that, by incorporating KG-CQR, the performance consistently outperforms the existing baseline in terms of retrieval effectiveness.
@@ 1 Introduction
These limitations become more pronounced when LLMs are applied to domain-specific knowledge retrieval, especially in addressing queries within vertical domains Bang et al. (2023). To address these challenges, recent research has explored the integration of Knowledge Graphs (KGs) into LLMs as a means to provide structured, accurate knowledge sources for enhanced reasoning Pan et al. (2024). KGs, which store facts in the form of triples (i.e., head entity, relation, tail entity), offer a robust and interpretable representation of knowledge. Consequently, there has been an increasing incorporation of KGs into applications based on LLMs to improve performance across various tasks, such as question answering Ding et al. (2024), fact verification Pham et al. (2025a), and recommendation systems Abu-Rasheed et al. (2024).
@@
This study focuses on enhancing the retrieval process for RAG systems by integrating KG technologies to enrich contextual information for complex input queries. Specifically, the objective is to tackle a critical challenge in current systems: the embedding misalignment between sentence-level queries and document-level corpus representations Ma et al. (2023). Accordingly, existing methods often employ large language models (LLMs) to decompose complex queries Mao et al. (2024) (Figure 1(a)). Nonetheless, in terms of retrieval performance, this approach frequently underperforms due to insufficient contextual alignment with the corpus. Subsequently, Gao et al. (2023) proposed a new approach by generating hypothetical documents to facilitate document-document similarity comparisons (Figure 1(b)). However, this method heavily relies on underlying LLMs, introducing risks of hallucination. In terms of knowledge-grounded expansion generation, Xia et al. (2025) introduced a knowledge-aware approach that leverages both unstructured data and structured relations. Nevertheless, their reliance on predefined relation schemas between entities (e.g., title) and documents constrains the scalability and adaptability.
@@ 2.1 Query Expansion using LLM
To handle complex queries effectively, query expansion is often essential for improving the performance of the retrieval process Azad and Deepak (2019). Traditional approaches decompose input queries into multi-view representations to enhance retrieval accuracy Zhang et al. (2022). Recently, with the rapid advancement of LLMs, a promising direction involves query enhancement, either through prompt-based techniques leveraging LLMs Wang et al. (2023), or by developing trainable frameworks that generate refined queries Mao et al. (2024). These methods aim to reformulate input queries to produce more effective semantic representations Chan et al. (2024); Chen et al. (2024). However, they still struggle to bridge the inherent gap between user queries and the knowledge corpus at the retrieval embedding level Liu et al. (2025). Accordingly, to further improve retrieval effectiveness, especially in domain-specific applications, a deeper exploitation of contextual generation remains essential Li et al. (2025).
@@ 2.2 Contextual Retrieval
Contextualized retrieval has emerged as an effective strategy for improving retrieval performance, particularly in complex and challenging settings Morris and Rush (2024). Recent methods, such as RAPTOR Sarthi et al. (2024), GraphRAG Edge et al. (2024), and HippoRAG Gutierrez et al. (2024), adopt recursive procedures that integrate embedding, clustering, and summarization techniques to construct hierarchical document representations using graph-based structures. Conceptually, these approaches follow a corpus-centric paradigm, wherein hierarchical structures are leveraged to enhance contextual retrieval across the original document corpus. In terms of query expansion through contextualization, Gao et al. (2023) proposes HyDE, a novel approach that leverages LLMs to generate hypothetical documents conditioned on the input query. Accordingly, the query is first processed by an LLM following specific instructions to produce hypothetical documents, which are then used as pseudo-contexts for retrieval based on document-to-document similarity. However, a key limitation of HyDE lies in its dependence on LLM-powered generated content, where potential inaccuracies or hallucinations can degrade retrieval effectiveness Zhang et al. (2024); Xia et al. (2025). Moreover, query expansion strategies must account for domain-specific context sensitivity, as the same entities may vary in meaning or relevance across different domains Bui et al. (2021). Therefore, this study proposes a novel contextual retrieval approach, which focuses on providing contextual information for the input query, based on the structured relation of the corpus-centric KG.
@@ 3.1.1 Structure Relation Representation
A corpus-centric KG includes a set of triplets (structure relation) $\mathcal{T}_{KG}$, which are defined as follows:
@@
Figure 2: Construction of structured relation representations using LLM-based prompting. Detailed prompt templates are provided in Appendix A.5.
@@
$TTR(\mathcal{T}_{KG}^{i})=llm(Promt_{ttr},D_{d}^{i},\mathcal{T}_{KG}^{i})$ (2)
@@ 3.2.1 Subgraph Extraction
$llm(Promt_{filter},q,\mathcal{T}_{KG}^{i})=True\}$
Here, $Prompt_{filter}$ denotes the instruction prompt used by the LLM for the final selection. The details of $Prompt_{filter}$ are provided in Appendix A.5.
@@ 3.2.2 Subgraph Completion
Notably, to reduce computational complexity in Step 2, instead of executing the naive BFS traversal, a limited number of nodes are expanded, guided by a heuristic function (BFSBeam). This function computes semantic similarity between the input query and aggregates the relevance scores of the TTRs along each path, which is illustrated in more detail in the Appendix A.3.
@@ 3.2.3 Contextual Generation
where $Prompt_{g}$ represents the generation instruction prompt, as detailed in Appendix A.5. The enriched subgraph $\hat{\mathcal{T^{\prime\prime}}}_{KG}$ serves as contextual input to the LLM, facilitating the generation of a contextually enriched query representation. This reformulated query can then be encoded within the same embedding space as the corpus documents, enabling effective retrieval.
@@ 4.1 Experimental Setup
RAGBench MultiHop-RAG Model mAP Recall@5 Recall@10 Recall@25 mAP Recall@5 Recall@10 Recall@25 BM25 0.329 0.337 0.399 0.462 0.241 0.261 0.353 0.486 DPR 0.276 0.286 0.348 0.425 0.099 0.125 0.183 0.284 BGE 0.521 0.51 0.589 0.657 0.227 0.251 0.357 0.52 QE + BM25 0.280 0.291 0.349 0.415 0.124 0.135 0.187 0.256 QE + DPR 0.251 0.260 0.317 0.392 0.058 0.069 0.101 0.169 QE + BGE 0.487 0.476 0.553 0.618 0.139 0.147 0.211 0.313 HyDE + DPR 0.286 0.293 0.354 0.426 0.106 0.127 0.188 0.297 HyDE + BGE 0.516 0.507 0.586 0.638 0.232 0.256 0.363 0.524 KG-CQR + BM25 0.398 0.398 0.454 0.514 0.250 0.267 0.372 0.532 KG-CQR + DPR 0.316 0.319 0.384 0.462 0.129 0.157 0.224 0.34 KG-CQR + BGE 0.542 0.529 0.61 0.675 0.24 0.261 0.371 0.525
Table 1: Retrieval performances on RAGBench and MultiHop-RAG datasets, using LLaMA-3.3-70B as the backbone
@@ 4.2 Main Results
Table 1 presents the evaluation results of the retrieval process on both datasets. Retrieval accuracy is evaluated using standard metrics such as mean Average Precision (mAP) and Recall@k, where $k\in\{5,10,25\}$. The reported results use $\alpha=0.7$ (Equation 9), which was found to yield the best performance (the selection of this value is further discussed in Appendix A.2.2). From the results, we draw the following observations:
i) Retrieval Performance: KG-CQR significantly improves retrieval performance across various retrieval backbones. On the RAGBench dataset, KG-CQR + BGE achieves the best performance overall, with an mAP of 0.542 and Recall@25 of 0.675, outperforming both the baseline models and the HyDE-enhanced variants. On the more challenging MultiHop-RAG dataset, KG-CQR + BM25 achieves the highest recall metrics (e.g., Recall@25 = 0.532), demonstrating KG-CQR’s capability to enhance retrieval accuracy over traditional methods.
ii) Impact of Query Expansion (QE): Compared with their respective baselines, QE-augmented models generally underperform across both datasets. For instance, QE + BM25 (mAP = 0.280 on RAGBench, 0.124 on MultiHop-RAG) performs notably worse than plain BM25, and similar degradations are observed for DPR and BGE backbones. This suggests that naive expansion of queries can introduce noise and semantic drift, which outweighs the potential benefits of richer lexical coverage. In contrast, KG-CQR achieves consistent improvements by leveraging structured knowledge for contextually grounded reformulations rather than unguided expansions.
iii) Contextual Accuracy: The comparatively lower performance of HyDE relative to its baselines indicates potential limitations in relying extensively on synthetic queries generated by Large Language Models (LLMs). Specifically, while HyDE offers a straightforward method for enhancing contextual understanding, its effectiveness is notably sensitive to the contextual reliability of the generated content. This highlights the constraints of inadequately grounded synthetic information in retrieval tasks.
@@ 4.3.1 Impact of LLM Backbone
RAGBench MultiHop-RAG Backbone mAP Recall@5 Recall@10 Recall@25 mAP Recall@5 Recall@10 Recall@25 LLaMA-3.2-3B 0.537 0.524 0.604 0.672 0.230 0.251 0.359 0.520 LLaMA-3.1-8B 0.538 0.526 0.606 0.672 0.235 0.255 0.370 0.522 LLaMA-3.3-70B 0.542 0.529 0.61 0.675 0.24 0.261 0.371 0.525
@@ 4.3.2 Ablation Study
Method Recall@5 Recall@10 Recall@25 W/O TTR 0.486 0.572 0.641 W/O Sub.Comp. 0.525 0.605 0.671 KG-CQR 0.529 0.61 0.675
Table 3: Ablation studies on main components of KG-CQR on RAGBench Dataset
As reported results, removing TTR (Equation 2) leads to the most pronounced drop in performance (e.g., Recall@25 decreases from 0.675 to 0.641), underscoring the importance of TTR in accurately extracting relevant subgraphs that preserve semantic alignment with the query. This confirms that converting structured KG information into textual form plays a critical role in aligning the knowledge with the retrieval task. Similarly, omitting the Subgraph Completion module also results in a notable performance degradation, though less severe than removing TTR. This suggests that while the initial subgraph extraction is vital, enriching the subgraph context via completion further improves the model’s ability to retrieve relevant documents.
@@ 4.3.3 Multi-Step Retrieval for RAG Task
Model Retrieval F1 $\uparrow$ Iter $\downarrow$ Score $\uparrow$ RAGBench LLaMA-3.2-3B 0.372 3.293 3.122 LLaMA-3.1-8B BM25 0.393 2.748 3.424 LLaMA-3.3-70B 0.431 1.912 3.449 LLaMA-3.2-3B KG-CQR +BM25 0.407 2.714 3.317 LLaMA-3.1-8B 0.410 1.834 3.603 LLaMA-3.3-70B 0.443 1.393 3.826 HotpotQA LLaMA-3.2-3B 0.613 2.586 3.828 LLaMA-3.1-8B BM25 0.662 2.591 3.955 LLaMA-3.3-70B 0.663 1.465 4.103 LLaMA-3.2-3B KG-CQR +BM25 0.648 2.450 4.106 LLaMA-3.1-8B 0.673 2.350 4.245 LLaMA-3.3-70B 0.700 1.280 4.278 MuSiQue LLaMA-3.2-3B 0.096 3.789 1.929 LLaMA-3.1-8B BM25 0.141 3.52 2.564 LLaMA-3.3-70B 0.374 2.042 3.206 LLaMA-3.2-3B KG-CQR +BM25 0.124 3.245 2.334 LLaMA-3.1-8B 0.223 3.089 2.783 LLaMA-3.3-70B 0.489 2.15 3.778
@@
To assess the generalizability of KG-CQR, experiments were conducted with three LLMs of varying sizes across multiple datasets. The evaluation highlights the role of KG-CQR in enhancing retrieval performance for reasoning-intensive RAG tasks. We randomly sampled 500 examples from the RAGBench test set and evaluated results using F1, GPT-Score ($Score$) Fu et al. (2024), and the average number of reasoning steps ($Iter$). In addition, two widely used multi-hop QA benchmarks, HotpotQA Yang et al. (2018) and MuSiQue Trivedi et al. (2022), were included in the evaluation. GPT-Score was computed using GPT-4o through the OpenAI API, based on its performance on the Judge LLM leaderboard [^2] https://huggingface.co/spaces/AtlaAI/judge-arena. As shown in Table 4, several key insights can be drawn: i) KG-CQR substantially improves retrieval quality across datasets: On RAGBench, KG-CQR + BM25 consistently outperforms BM25, with performance gains across all LLM sizes (e.g., F1 improves from 0.393 to 0.410 on LLaMA-3.1-8B). Similar improvements are observed on HotpotQA, where KG-CQR yields a significant gain for the largest model (F1 = 0.700 vs. 0.663). The effect is most pronounced on MuSiQue, where KG-CQR + BM25 achieves F1 = 0.489 with LLaMA-3.3-70B compared to 0.374 for BM25, underscoring its effectiveness for complex multi-hop reasoning (results with the BGE retriever are provided in Appendix A.2.1). ii) Contextualized reformulations reduce reasoning iterations: KG-CQR consistently decreases the average number of reasoning steps. For example, on HotpotQA with LLaMA-3.3-70B, the number of steps is reduced from 1.465 to 1.280. This suggests that knowledge-grounded query reformulations provide more accurate intermediate evidence, enabling models to converge on answers with fewer redundant reasoning cycles. iii) Cross-model scalability and robustness: Performance gains are observed across different LLM sizes, highlighting the adaptability of KG-CQR. Notably, the improvements are more pronounced on datasets requiring deeper reasoning (e.g., RAGBench and MuSiQue), indicating that KG-CQR effectively complements LLM reasoning by supplying better-targeted retrieval contexts.

### v3

@@ Abstract
The integration of knowledge graphs (KGs) with large language models (LLMs) offers significant potential to enhance the retrieval stage in retrieval-augmented generation (RAG) systems. In this study, we propose KG-CQR [^1] https://github.com/tnmai59/KG-CQR, a novel framework for Contextual Query Retrieval (CQR) that enhances the retrieval phase by enriching complex input queries with contextual representations derived from a corpus-centric KG. Unlike existing methods that primarily address corpus-level context loss, KG-CQR focuses on query enrichment through structured relation representations, extracting and completing relevant KG subgraphs to generate semantically rich query contexts. Comprising subgraph extraction, completion, and contextual generation modules, KG-CQR operates as a model-agnostic pipeline, ensuring scalability across LLMs of varying sizes without additional training. Experimental results on the RAGBench and MultiHop-RAG datasets demonstrate that KG-CQR outperforms strong baselines, achieving improvements of up to 4–6% in mAP and approximately 2–3% in Recall@25. Furthermore, evaluations on challenging RAG tasks such as multi-hop question answering show that, by incorporating KG-CQR, the performance outperforms the existing baseline in terms of retrieval effectiveness.
@@ 1 Introduction
These limitations are exacerbated when LLMs are applied to domain-specific knowledge retrieval, especially in addressing queries within vertical domains Bang et al. (2023). To address these challenges, recent research has explored the integration of knowledge graphs (KGs) into LLMs as a means to provide structured, accurate knowledge sources for enhanced reasoning Pan et al. (2024). KGs, which store facts in the form of triples (i.e., head entity, relation, tail entity), offer a robust and interpretable representation of knowledge. Consequently, KGs have been increasingly incorporated into applications based on LLMs to improve performance across various tasks, such as question answering Ding et al. (2024), fact verification Pham et al. (2025a), and recommendation systems Abu-Rasheed et al. (2024).
@@
This study focuses on enhancing the retrieval process for RAG systems by integrating KG technologies to enrich contextual information for complex input queries. Specifically, the objective is to tackle a critical challenge in current systems: misalignment between query and document embeddings Ma et al. (2023). Accordingly, existing methods often employ LLMs to decompose complex queries Mao et al. (2024) (Figure 1(a)). Nonetheless, in terms of retrieval performance, this approach frequently underperforms due to insufficient contextual alignment with the corpus. Subsequently, Gao et al. (2023) proposed a new approach by generating hypothetical documents to facilitate document-document similarity comparisons (Figure 1(b)). However, this method heavily relies on underlying LLMs, introducing risks of hallucination. In terms of knowledge-grounded expansion generation, Xia et al. (2025) introduced a knowledge-aware approach that leverages both unstructured data and structured relations. Nevertheless, their reliance on predefined relation schemas between entities (e.g., title) and documents constrains the scalability and adaptability.
@@ 2.1 Query Expansion using LLM
To handle complex queries effectively, query expansion is often essential for improving the performance of the retrieval process Azad and Deepak (2019). Traditional approaches decompose input queries into multi-view representations to enhance retrieval accuracy Zhang et al. (2022). Recently, with the rapid advancement of LLMs, a promising direction involves query enhancement, either through prompt-based techniques leveraging LLMs Wang et al. (2023), or by developing trainable frameworks that generate refined queries Mao et al. (2024). These methods aim to reformulate queries into more effective semantic representations Chan et al. (2024); Chen et al. (2024). However, they still struggle to bridge the inherent gap between queries and the knowledge corpus within the retrieval embedding space Liu et al. (2025). Accordingly, to further improve retrieval effectiveness, especially in domain-specific applications, a deeper exploitation of contextual generation remains essential Li et al. (2025).
@@ 2.2 Contextual Retrieval
Contextualized retrieval has emerged as an effective strategy for improving retrieval performance, particularly in complex and challenging settings Morris and Rush (2024). Recent methods, such as RAPTOR Sarthi et al. (2024), GraphRAG Edge et al. (2024), and HippoRAG Gutierrez et al. (2024), adopt recursive procedures that integrate embedding, clustering, and summarization techniques to construct hierarchical document representations using graph-based structures. Conceptually, these approaches follow a corpus-centric paradigm, wherein hierarchical structures are leveraged to enhance contextual retrieval across the original document corpus. In terms of query expansion through contextualization, Gao et al. (2023) proposes HyDE, a novel approach that leverages LLMs to generate hypothetical documents conditioned on the input query. Accordingly, the query is first processed by an LLM following specific instructions to produce hypothetical documents, which are then used as pseudo-contexts for retrieval based on document-to-document similarity. However, a key limitation of HyDE lies in its dependence on LLM-generated content, where potential inaccuracies or hallucinations can degrade retrieval effectiveness Zhang et al. (2024); Xia et al. (2025). Moreover, query expansion strategies must account for domain-specific context sensitivity, as the same entities may vary in meaning or relevance across different domains Bui et al. (2021). Therefore, this study proposes a novel contextual retrieval approach, which focuses on providing contextual information for the input query, based on the structured relation of the corpus-centric KG.
@@ 3.1.1 Structure Relation Representation
A corpus-centric KG includes a set of triplets (structured relations) $\mathcal{T}_{KG}$, which are defined as follows:
@@
Figure 2: Construction of structured relation representations using LLM-based prompting. Detailed prompt templates are provided in Appendix 6.5.
@@
$TTR(\mathcal{T}_{KG}^{i})=llm(Prompt_{ttr},D_{d}^{i},\mathcal{T}_{KG}^{i})$ (2)
@@ 3.2.1 Subgraph Extraction
$llm(Prompt_{filter},q,\mathcal{T}_{KG}^{i})=True\}$
Here, $Prompt_{filter}$ denotes the instruction prompt used by the LLM for the final selection. The details of $Prompt_{filter}$ are provided in Appendix 6.5.
@@ 3.2.2 Subgraph Completion
Notably, to reduce computational complexity in Step 2, instead of executing the naive BFS traversal, a limited number of nodes are expanded, guided by a heuristic function (BFSBeam). This function computes semantic similarity between the input query and aggregates the relevance scores of the TTRs along each path, which is illustrated in more detail in the Appendix 6.3.
@@ 3.2.3 Contextual Generation
where $Prompt_{g}$ represents the generation instruction prompt, as detailed in Appendix 6.5. The enriched subgraph $\hat{\mathcal{T^{\prime\prime}}}_{KG}$ serves as contextual input to the LLM, facilitating the generation of a contextually enriched query representation. This reformulated query can then be encoded within the same embedding space as the corpus documents, enabling effective retrieval.
@@ 4.1 Experimental Setup
RAGBench MultiHop-RAG Model mAP Recall@5 Recall@10 Recall@25 mAP Recall@5 Recall@10 Recall@25 BM25 0.329 0.337 0.399 0.462 0.241 0.261 0.353 0.486 DPR 0.276 0.286 0.348 0.425 0.099 0.125 0.183 0.284 BGE 0.521 0.510 0.589 0.657 0.227 0.251 0.357 0.520 QE + BM25 0.280 0.291 0.349 0.415 0.124 0.135 0.187 0.256 QE + DPR 0.251 0.260 0.317 0.392 0.058 0.069 0.101 0.169 QE + BGE 0.487 0.476 0.553 0.618 0.139 0.147 0.211 0.313 HyDE + DPR 0.286 0.293 0.354 0.426 0.106 0.127 0.188 0.297 HyDE + BGE 0.516 0.507 0.586 0.638 0.232 0.256 0.363 0.524 KG-CQR + BM25 0.398 0.398 0.454 0.514 0.250 0.267 0.372 0.532 KG-CQR + DPR 0.316 0.319 0.384 0.462 0.129 0.157 0.224 0.340 KG-CQR + BGE 0.542 0.529 0.610 0.675 0.240 0.261 0.371 0.525
Table 1: Retrieval performance on the RAGBench and MultiHop-RAG datasets with LLaMA-3.3-70B as the backbone LLM for contextual generation
@@ 4.2 Main Results
Table 1 presents the evaluation results of the retrieval process on both datasets. Retrieval accuracy is evaluated using standard metrics such as mean Average Precision (mAP) and Recall@k, where $k\in\{5,10,25\}$. The reported results use $\alpha=0.7$ (Equation 9), which was found to yield the best performance (the selection of this value is further discussed in Appendix 6.2.2). From the results, we draw the following observations:
i) Retrieval Performance: KG-CQR significantly improves retrieval performance across various retrieval backbones. On the RAGBench dataset, KG-CQR + BGE achieves the best performance overall, with an mAP of 0.542 and Recall@25 of 0.675, outperforming both the baseline models and the HyDE-enhanced variants. On the more challenging MultiHop-RAG dataset, KG-CQR + BM25 achieves the highest recall metrics (e.g., Recall@25 = 0.532), demonstrating KG-CQR’s effectiveness compared to traditional methods.
ii) Impact of Query Expansion (QE): Compared with their respective baselines, QE-augmented models generally underperform across both datasets. For instance, QE + BM25 (mAP = 0.280 on RAGBench, 0.124 on MultiHop-RAG) performs notably worse than plain BM25, and similar degradations are observed for DPR and BGE backbones. This suggests that naive query expansion often introduces noise and semantic drift, which outweighs the potential benefits of richer lexical coverage. In contrast, KG-CQR achieves consistent improvements by leveraging structured knowledge for contextually grounded reformulations instead of unguided expansions.
iii) Contextual Accuracy: The comparatively lower performance of HyDE relative to its baselines indicates potential limitations in relying extensively on synthetic queries generated by LLMs. Specifically, while HyDE offers a straightforward method for enhancing contextual understanding, its effectiveness is notably sensitive to the contextual reliability of the generated content. This highlights the constraints of inadequately grounded synthetic information in retrieval tasks.
@@ 4.3.1 Impact of LLM Backbone
RAGBench MultiHop-RAG Backbone mAP Recall@5 Recall@10 Recall@25 mAP Recall@5 Recall@10 Recall@25 LLaMA-3.2-3B 0.537 0.524 0.604 0.672 0.230 0.251 0.359 0.520 LLaMA-3.1-8B 0.538 0.526 0.606 0.672 0.235 0.255 0.370 0.522 LLaMA-3.3-70B 0.542 0.529 0.610 0.675 0.240 0.261 0.371 0.525
@@ 4.3.2 Ablation Study
Method Recall@5 Recall@10 Recall@25 W/O TTR 0.486 0.572 0.641 W/O Sub.Comp. 0.525 0.605 0.671 KG-CQR 0.529 0.610 0.675
Table 3: Ablation study of KG-CQR components on RAGBench
As shown in the results, removing TTR (Equation 2) leads to the most pronounced drop in performance (e.g., Recall@25 decreases from 0.675 to 0.641), underscoring the importance of TTR in accurately extracting relevant subgraphs that preserve semantic alignment with the query. This confirms that converting structured KG information into textual form plays a critical role in aligning the knowledge with the retrieval task. Similarly, omitting the Subgraph Completion module also results in a notable performance degradation, though less severe than removing TTR. This suggests that while the initial subgraph extraction is vital, enriching the subgraph context via completion further improves the model’s ability to retrieve relevant documents.
@@ 4.3.3 Multi-Step Retrieval for RAG Task
Model Retrieval F1 $\uparrow$ Iter $\downarrow$ Score $\uparrow$ RAGBench LLaMA-3.2-3B 0.372 3.293 3.122 LLaMA-3.1-8B BM25 0.393 2.748 3.424 LLaMA-3.3-70B 0.431 1.912 3.449 LLaMA-3.2-3B KG-CQR +BM25 0.407 2.714 3.317 LLaMA-3.1-8B 0.410 1.834 3.603 LLaMA-3.3-70B 0.443 1.393 3.826 HotpotQA LLaMA-3.2-3B 0.613 2.586 3.828 LLaMA-3.1-8B BM25 0.662 2.591 3.955 LLaMA-3.3-70B 0.663 1.465 4.103 LLaMA-3.2-3B KG-CQR +BM25 0.648 2.450 4.106 LLaMA-3.1-8B 0.673 2.350 4.245 LLaMA-3.3-70B 0.700 1.280 4.278 MuSiQue LLaMA-3.2-3B 0.096 3.789 1.929 LLaMA-3.1-8B BM25 0.141 3.52 2.564 LLaMA-3.3-70B 0.374 2.042 3.206 LLaMA-3.2-3B KG-CQR +BM25 0.124 3.245 2.334 LLaMA-3.1-8B 0.223 3.089 2.783 LLaMA-3.3-70B 0.489 2.150 3.778
@@
To assess the generalizability of KG-CQR, experiments were conducted with three LLMs of varying sizes across multiple datasets. The evaluation highlights the role of KG-CQR in enhancing retrieval performance for reasoning-intensive RAG tasks. We randomly sampled 500 examples from the RAGBench test set and evaluated results using F1, GPT-Score ($Score$) Fu et al. (2024), and the average number of reasoning steps ($Iter$). In addition, two widely used multi-hop QA benchmarks, HotpotQA Yang et al. (2018) and MuSiQue Trivedi et al. (2022), were included in the evaluation. GPT-Score was computed using GPT-4o through the OpenAI API, based on its performance on the Judge LLM leaderboard [^2] https://huggingface.co/spaces/AtlaAI/judge-arena. As shown in Table 4, several key insights can be drawn: i) KG-CQR substantially improves retrieval quality across datasets: On RAGBench, KG-CQR + BM25 consistently outperforms BM25, with performance gains across all LLM sizes (e.g., F1 improves from 0.393 to 0.410 on LLaMA-3.1-8B). Similar improvements are observed on HotpotQA, where KG-CQR yields a significant gain for the largest model (F1 = 0.700 vs. 0.663). The effect is most pronounced on MuSiQue, where KG-CQR + BM25 achieves F1 = 0.489 with LLaMA-3.3-70B compared to 0.374 for BM25, underscoring its effectiveness for complex multi-hop reasoning (results with the BGE retriever are provided in Appendix 6.2.1). ii) Contextualized reformulations reduce reasoning iterations: KG-CQR consistently decreases the average number of reasoning steps. For example, on HotpotQA with LLaMA-3.3-70B, the number of steps is reduced from 1.465 to 1.280. This suggests that knowledge-grounded query reformulations provide more accurate intermediate evidence, enabling models to converge on answers with fewer redundant reasoning cycles. iii) Cross-model scalability and robustness: Performance gains are observed across different LLM sizes, highlighting the adaptability of KG-CQR. Notably, the improvements are more pronounced on datasets requiring deeper reasoning (e.g., RAGBench and MuSiQue), indicating that KG-CQR effectively complements LLM reasoning by supplying better-targeted retrieval contexts.